# Revisiting the Jaynes–Cummings model with time-dependent coupling

### v1

@@ Acknowledgments
We thank Dr. Alison A. Silva for helpful discussions. This work was financially supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES, Finance Code 001), Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Instituto Nacional de Ciência e Tecnologia de Informação Quântica (INCT-IQ), Brazil, and the Air Force Office of Scientific Research (AFOSR), USA. DC would like to acknowledge financial support from Instituto Serrapilheira, and the Pró-Reitoria de Pesquisa e Inovação (PRPI) from the Universidade de São Paulo (USP) by financial support through the Programa de Estímulo à Supervisão de Pós-Doutorandos por Jovens Pesquisadores. F.M.A. acknowledges the CNPq Grant No 313124/2023-0, and A.V.-B. acknowledges the AFOSR award No FA9550-24-1-0009.

### v2

@@ Acknowledgments
We thank Dr. Alison A. Silva for helpful discussions. This work was financially supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES, Finance Code 001), Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Instituto Nacional de Ciência e Tecnologia de Informação Quântica (INCT-IQ), Brazil, and the Air Force Office of Scientific Research (AFOSR), USA. D.C. would like to acknowledge financial support from Instituto Serrapilheira, and the Pró-Reitoria de Pesquisa e Inovação (PRPI) from the Universidade de São Paulo (USP) by financial support through the Programa de Estímulo à Supervisão de Pós-Doutorandos por Jovens Pesquisadores. F.M.A. acknowledges the CNPq Grant No 313124/2023-0, F.M.A. and A.S.M.C. acknowledge Fundação Araucária Project No 305, and A.V.-B. acknowledges the AFOSR award No FA9550-24-1-0009.
Funding This work was supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES, Finance Code 001), the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), the Instituto Nacional de Ciência e Tecnologia de Informação Quântica (INCT-IQ), Brazil, and the Air Force Office of Scientific Research (AFOSR), USA. D.C. was supported by Instituto Serrapilheira and by the Pró-Reitoria de Pesquisa e Inovação (PRPI) from the Universidade de São Paulo (USP) through the Programa de Estímulo à Supervisão de Pós-Doutorandos por Jovens Pesquisadores. F.M.A. acknowledges support from CNPq (Grant No. 313124/2023-0) and Fundação Araucária (Project No. 305). A.S.M.C. acknowledges support from Fundação Araucária (Project No. 305) A.V.-B. acknowledges support from AFOSR (Award No. FA9550-24-1-0009).